a light year is a measure of lenth(not time)equal to distance that light in 1year.compute the conversion factors betweenlight year and meter. find distance to the star proxima centauri(4.0*10 to the power of 16m )in light years.


 


 


 


 


 


 


 


 


 


 


 

Grade:

1 Answers

vijesh jayan
12 years ago

365 *24*3600*300000*1000 = 9,460,800,000,000,000 m
= 9.46*1015 m

1 light year = 9.46*1015 m

(4*1016)/(9.46*1015)*1 light year = 4*1016 m